In a plant box consisting of 120 seedlings, 80 seedlings were treated with growth enhancer. Estimate p and q. What is the Point Estimate and Standard Deviation?

1
Expert's answer
2022-04-29T10:00:21-0400
"p=\dfrac{80}{120}=\dfrac{2}{3}"

"q=1-p=1-\dfrac{2}{3}=\dfrac{1}{3}"

Point Estimate


"\mu=np=120(\dfrac{2}{3})=80"

Standard Deviation


"\sigma=\sqrt{npq}=\sqrt{120(\dfrac{2}{3})(\dfrac{1}{3})}=\dfrac{4\sqrt{15}}{3}"